### v1

# Configuration Interaction Guided Sampling with Interpretable Restricted Boltzmann Machine

Abstract
We propose a data-driven approach using a Restricted Boltzmann Machine (RBM) to solve the Schrödinger equation in configuration space. Traditional Configuration Interaction (CI) methods, while powerful, are computationally expensive due to the large number of determinants required. Our approach leverages RBMs to efficiently identify and sample the most significant determinants, accelerating convergence and reducing computational cost. This method achieves up to 99.99% of the correlation energy even by four orders of magnitude less determinants compared to full CI calculations and up to two orders of magnitude less than previous state of the art works. Additionally, our study demonstrate that the RBM can learn the underlying quantum properties, providing more detail insights than other methods . This innovative data-driven approach offers a promising tool for quantum chemistry, enhancing both efficiency and understanding of complex systems.
keywords:
Full Configuration Interaction (FCI); Restricted Boltzmann Machine (RBM); Data-Driven Quantum Chemistry.
journal: Computer Physics Communications
1 Introduction
Understanding the quantum properties of molecules and materials is crucial in various scientific and technological fields. The Schrödinger equation serves as the fundamental base for this understanding, but its exact solution is often intractable. The Configuration Interaction (CI) method offers a powerful approximation, but it suffers from the fact that the vast number of determinants are required for realistic systems. This complexity limits the applicability of traditional CI methods to many important molecules.
This work proposes a novel approach that harnesses the power of artificial intelligence, specifically Restricted Boltzmann Machines (RBMs), to address the limitations of CI. RBMs are probabilistic neural networks that are experts at efficiently sampling prohibitive spaces to find the current distribution. In recent years have been used in the field of Computational Biology Medicine and Chemistry in pattern extraction and learning complex distributions [1, 2, 3].
Although new generative models have been developed in artificial intelligence such as Generative Adversarial Networks (GAN), Variational Autoencoders(VAE) or Deep learners, RBMs have become popular because they only have one visible and one hidden layer, resulting in simple and potentially interpretable learned features.
Our strategy uses the fact that, like Principal Component Analysis (PCA), an RBM can learn the principal components that maximize the variance in a dataset. This allows us to obtain the main characteristics. However, unlike PCA, which has linear transformations of the data vectors, RBMS are non-linear in nature and have the ability to learn more complicated relationships between visible and hidden units. In this work, we use an RBM to identify the most relevant determinants in the configuration space, accelerating convergence in solving the Schrödinger equation compared to random sampling or traditional CI methods.
The determinants, are fundamental to CI methods, such determinants are governed by Slater’s rules [4]. These rules dictate that only determinants differing by one or two permutations have non-zero Hamiltonian matrix elements. This implies an inherent relationship among the determinants, suggesting that those making the most significant contributions to the wavefunction share common features. Our RBM-based approach exploits this relationship by learning the underlying distribution of these determinants that contribute more to the solution of the Schrödinger equation.
It is noteworthy that the CI method can be approximated using various levels of theory, such as configuration interaction with single and double excitations (CISD) and considering triple, and quadruple excitations too (CISDTQ). CISD, for instance, can recover approximately $95\%$ of the correlation energy (in small molecules at equilibrium geometries) [5, 6], while CISDTQ can achieve up to $99.8\%$ of the full CI energy [7]. These approximations highlight the importance of employing higher-order methods like Full Configuration Interaction (FCI) to obtain more accurate results.
Traditional approaches to CI, such as Quantum Monte Carlo or stochastic search methods [8], can be computationally inefficient. In contrast, RBMs, while incorporating a stochastic process at their core (Gibbs Sampling or Contrastive Divergence), ultimately guides the sampling process based on the learned data distribution. This guided approach enhances the efficiency and effectiveness of the RBM-based CI method.
Previous works has explored the use of RBMs to solve the CI approximation [9, 10]. In this study, we present a significant improvement to the state of the art by introducing an innovative and differentiated model training methodology.
Specifically, our approach utilizes RBM as generative model with a innovative guided model strategy that focuses on iteratively selecting determinants with higher probabilities, thereby facilitating faster and more precise solution convergence.
A key result of this study is the ability to reach similar or even better results as current state of the art methods, with up to two orders of magnitude less determinants. Such improvement is obtained by implementing a taboo list that allows for retaining only the determinants that contribute the most to the final wavefunction in each iteration, and using that information for subsequent training when generating new determinants while searching the Hilbert space.
This strategy helps us maintain the time complexity and prevent rapid growth in computational costs, reducing overall computation time.
Furthermore, we present, a novel guided learning strategy for the RBM model, and demonstrate that such model is explainable. The insights learned from the RBM model provide a deeper understanding of the underlying patterns that make up the solution. This investigation revealed that the model appears to discern a distinctive pattern across various molecules, which remains consistent within the same molecule using different molecular basis. This pattern seems linked to the radial distribution of probability of specific molecular bonds and the energies of energy orbitals. This suggests that the model has uncovered a previously unrecognized relationship among determinants. Such interpretability not only enhances current applications but also lays the foundation for future research in applying artificial intelligence techniques to quantum chemistry.
In summary, this study represents a significant advance in the field of computational quantum chemistry by introducing an improved CI solution strategy, supported by a guided RBM learning model, accompanied by a practical implementation and a detailed analysis of the results.
The paper begins with Section 2, which provides a theoretical background on solving the Schrödinger equation for complex systems (2.1). It also covers the fundamentals of the Configuration Interaction method and its limitations (2.2), along with an overview of Restricted Boltzmann Machines (2.3).
In Subsection (2.4), we introduce our proposed model that utilizes an RBM as a generative model, highlighting key features and innovations aimed at approximating FCI. We then present the results and discussions (3), followed by the conclusions in Section (4).
2 Material and methods
2.1 Many-body Hamiltonian
In quantum mechanics, solving the time-independent Schrödinger equation (Eq. 1) is the cornerstone for predicting the behavior of a system:
$H\psi=E\psi.$ (1)
Here, $H$ represents the system’s Hamiltonian, $\psi$ is the wave function, and $E$ is the corresponding energy eigenvalue. Solving this eigenvalue equation provides the complete energy spectrum and wave functions, enabling the calculation of all properties and predictions for the system.
However, Eq. 1 presents a significant challenge when dealing with many-body problems. Consider a solid-state system with a vast number of microscopic constituents (atoms) and constituent particles (electrons). An accurate description requires a Hamiltonian that incorporates the kinetic energies of both electrons and nuclei, along with all interaction terms: electron-electron, electron-nuclei, and nuclei-nuclei. This leads to a more comprehensive form of the Schrödinger eigenvalue equation given by Eq. 2:
$\left[\frac{h^{2}}{2m}\sum_{i=1}^{N}\nabla_{i}^{2}+\sum_{i=1}^{N}V(r_{i})+\right.\left.\sum_{i=1}^{N}\sum_{j<1}U(r_{i},r_{j})\right]\psi=E\psi.$ (2)
Here, $m$ represents the electron mass. The first term describes the kinetic energy of each electron due to their individual momenta. The second term, $V(r_{i})$, represents the potential energy of each electron interacting with the collection of atomic nuclei, often referred to as the external potential. This term can be further expanded as:
$\sum_{i=1}^{N}V(r_{i})=\sum_{i=1}^{N}\sum_{j=1}^{N_{atoms}}\frac{-Z_{j}}{|r_{i}-R_{j}|},$ (3)
where $Z_{j}$ is the atomic number (number of protons) of the $j$-th nucleus, and $r_{i}$ and $R_{A}$ are the position vectors of the $i$-th electron and $j$-th nucleus, respectively. The negative sign indicates the attractive nature of the interaction between negatively charged electrons and positively charged nuclei.
Considering the Born-Oppenheimer approximation, the nuclei are considered stationary due to their much larger mass compared to electrons. Therefore, they exert a fixed Coulombic potential on the electrons [11].
The third term in the Hamiltonian, $U(r_{i},r_{j}$), represents the potential energy arising from the interaction between different electrons due to their Coulombic repulsion. This term is expressed as:
$\sum_{i=1}^{N}\sum_{j<1}U(r_{i},r_{j})=\sum_{i=1}^{N}\sum_{j<1}\frac{e^{2}}{2|r_{i}-r_{j}|}.$ (4)
In Eq. 2, $\psi$ represents the electronic wave function, a function that depends on the spatial coordinates of all $N$ electrons: $\psi=\psi(r_{1},...r_{n})$. The solutions of the Schrödinger equation with this Hamiltonian provide the corresponding energy levels ($E$) of the electronic system.
The complexity arises due to the high dimensionality of the electronic wave function. For instance, a single CO2 molecule requires a 66-dimensional wave function (3 spatial coordinates for each of the 22 electrons). For a larger system, like a nanocluster with 100 Pt atoms, the wave function becomes a staggering 23,000-dimensional object. This immense complexity makes it computationally intractable to solve the Schrödinger equation exactly for most many-body systems [12].
2.1.1 Challenges and Approximation Methods:
Due to the high dimensionality of the many-body problem, various approximation methods are employed to solve the Schrödinger equation. Two of the most well-established techniques are wave function-based methods and density-based methods. Wave function-based methods include perturbative and variational approaches, such as the Configuration Interaction (CI) method, which will be the primary focus of the following section. CI explicitly considers a linear combination of multiple electronic configurations to improve the accuracy of the wave function.
On the other hand, Density Functional Theory (DFT) is a popular density-based method known for its computational efficiency. However, DFT’s application can still suffer from large pervasive errors that cause qualitative failures in predicted properties. For instance, DFT often underestimates the barriers of chemical reactions, the band gaps of materials, the energies of dissociating molecular ions, and charge transfer excitation energies. Additionally, it tends to overestimate the binding energies of charge transfer complexes and the response to an electric field in molecules and materials [13].
In contrast, CI, particularly Full Configuration Interaction (FCI), offers a rigorous wave function-based approach that does not rely on approximations of the exchange-correlation functional. This makes CI a preferred choice in cases where the highest accuracy is required [14], such as in the study of bond-breaking processes, excited states, and systems with near-degenerate electronic states. Despite its higher computational cost, CI provides a more detailed and accurate description of electron correlation effects.
We will delve deeper into the details and applications of CI, specifically FCI, in the next section.
2.2 Configuration Interaction
There are well-established methods for approximating the wave function in order to determine the ground state of electronic systems. One prominent method is the Hartree-Fock method, which simplifies the problem by assuming that each electron moves independently in a central potential created by the other electrons. This eliminates the need to consider the $r_{i}$ term in the many-body Hamiltonian equation. However, the accuracy of the ground state energies obtained using this method is limited due to the use of an average potential to account for electron-electron interactions [6].
To overcome this limitation, Configuration Interaction (CI) methods are employed. These methods acknowledge that electrons interact and repel each other, leading to correlated motion within atoms. Instead of being restricted to a single orbital configuration represented by a Slater determinant, CI methods construct properly antisymmetrized product wave functions using multiple configurations [15]. By distributing electrons among different orbitals, spatial correlations in electron motion can be taken into account. CI utilizes a variational wave function [16] mathematically expressed as shown in Eq. 5:
$\Psi=\sum_{i=0}^{I}c_{i}\Phi_{i}=c_{0}\Phi_{0}+c_{1}\Phi_{1}+\ldots$ (5)
In this expansion of the total wave function ($\Psi$), it is expressed as a linear combination of diferent electron configurations ($\Phi_{i}$).The bases used in the expansion are usually excited Slater determinants [17] . The expansion coefficients ($c_{i}$) determine the contribution of each configuration to the total wave function.
The Hamiltonian matrix is constructed so that $H_{ij}=\langle\Phi_{i}|\hat{H}|\Phi_{j}\rangle$ for $i,j=0,1,2,\ldots,I$. Using Slater’s rules allows us to express the matrix elements $H_{ij}$ in terms of one- and two-electron integrals. The one-electron integrals are written as [15]:
[…]
and the two-electron integrals are written as:
[…]
Here, $\phi_{i}$ and $\phi_{j}$ are spin-orbitals, which are functions of both spatial and spin coordinates of a single electron. The Hamiltonian is then diagonalized to obtain the coefficients $c_{i}$ in the expansion of the wave function.
When the expansion in Eq. 5 includes all possible electron configurations with the appropriate symmetry, it constitutes a full configuration interaction (FCI) procedure that exactly solves the electronic Schrödinger equation within the space spanned by the one-particle basis set.
If only determinants of single and double excitations are considered, the method is known as Configuration Interaction Singles Doubles (CISD) [18]. All combinations, or every possible determinant is referred to as Full Configuration Interaction (FCI), providing the most accurate approximation to the ground state within the specific basis set. However, the FCI approach as a complete solution, suffers from poor scalability as the number of determinants grows exponentially with $\binom{2K}{N}$ [18], where $2K$ is the number of spin-orbitals and $N$ is the number of particles [10].
The total number of Slater determinants, $N_{det}$, is calculated as follows (assuming a closed-shell singlet state):
$N_{det}=\left(\frac{N_{o}!}{(N_{o}-N_{e})!N_{e}!}\right)^{2}\approx\binom{N_{o}}{N_{e}/2}^{2}$ (8)
where $N_{o}$ represents the number of molecular orbitals (MOs) and $N_{e}=N_{\alpha}=N_{\beta}$ represents the number of electrons with $\alpha$-spin, which is half of the total number of electrons [19]. This formula highlights the rapid growth in the number of determinants as the size of the system increases, contributing to the computational challenges of the FCI approach.
The following Table 1 shows the number of determinants required for simple molecules, demonstrating that a significant number of determinants are needed even for these simple systems.
System Basis Active Space FCI $N_{dets}$
[…]
Table 1: Required number of determinants for each molecule using different basis sets in FCI calculations
This exponential growth in the number of determinants with system size renders FCI intractable for large molecules or complex materials. While FCI provides the most accurate solution within a chosen basis set, exploring this vast configuration space becomes computationally intractable.
To overcome this prohibitive configuration space challenge, recent advancements in the field of machine learning offer promising avenues. Techniques like Restricted Boltzmann Machines (RBMs) can be employed as generative models [20, 21]. These models can learn the underlying probability distribution of the electronic system and efficiently sample the most relevant configurations for the FCI calculation. This selective approach, guided by the RBM, significantly reduces the computational cost compared to traditional methods that explore the configuration space randomly. The next section will delve into how RBMs can be harnessed to address the limitations of FCI and pave the way for studying electronic structures in complex systems.
2.3 Restricted Boltzmann Machine
RBM (Restricted Boltzmann Machines) is a type of generative model and unsupervised learning algorithm used in the field of machine learning. RBM is a probabilistic energy-based neural network consisting of two layers of nodes: a visible layer and a hidden layer. The nodes in the visible layer represent observed variables, while the nodes in the hidden layer represent latent or unobserved features. These two layers are fully connected to each other, but there are no connections within each layer (See Figure 1).
Figure 1: General architecture of the Restricted Boltzmann Machines (RBM).
The main goal of an RBM is to learn the synaptic weights that maximize the joint probability of the observed data in the visible layer and the latent representations in the hidden layer, or in other words the RBM as a generative model, represents a probability distribution (over states of the visible units) where low-energy configurations have higher probability. The energy is determined by the connection weights which are the parameters to be learnt from the data to built a generative model that learns the underlaying dynamic of the system [22]. This is achieved using an algorithm called “Gibbs sampling” or “Contrastive Divergence” [23].
The learning process of an RBM involves initializing the synaptic weights randomly and propagating the data forward and backward in the network to calculate activations and reconstructions. The learning algorithm adjusts the synaptic weights to maximize the probability that the observed data is effectively reconstructed in the visible layer.
A RBM can be mathematically described using the energy of a network and the associated probability distribution. The energy of an RBM is defined in terms of the activation states of the visible and hidden nodes, as well as the synaptic weights connecting these layers. The energy is denoted as $E(v,h)$, where $v$ represents the vector of states of the visible nodes and $h$ represents the vector of states of the hidden nodes.
The energy of an RBM is defined as follows:
$E(v,h)=-\sum_{i}a_{i}v_{i}-\sum_{j}b_{j}h_{j}-\sum_{i,j}v_{i}h_{j}w_{ij}.$
Where:
  • 1.
    $a_{i}$ and $b_{j}$ are the biases of the visible and hidden nodes, respectively.
  • 2.
    $w_{ij}$ is the synaptic weight between visible node $i$ and hidden node $j$.
  • 3.
    $\sum_{i}a_{i}v_{i}$ and $\sum_{j}b_{j}h_{j}$ represent the contributions of the biases.
  • 4.
    $\sum_{i,j}v_{i}h_{j}w_{ij}$ represents the interactions between the visible and hidden nodes.
The network assigns a probability to every possible pair of a visible and a hidden vector via the energy function shown in Eq. 9:
$p(v,h)=\frac{1}{Z}e^{-E(v,h)}.$ (9)
This joint probability $p(v,h)$ indicates the probability that both $v$ and $h$ are in a specific state.
The probability distribution associated with an RBM is defined using the partition function $Z$, which normalizes the energy over all possible visible and hidden states:
$Z=\sum_{v}\sum_{h}e^{-E(v,h)},$
where the summations are performed over all possible visible and hidden states.
Therefore, the probability assigned by the network to a visible vector $v$ is determined by summing over all possible hidden vectors $h$:
$P(v)=\frac{1}{Z}\sum_{h}\exp(-E(v,h)).$ (10)
By adjusting the weights and biases, the network can increase the probability assigned to a training input. This adjustment lowers the energy of the target input while raising the energy of other inputs, particularly those with lower energies that significantly contribute to the partition function.
Therefore, the parameters that maximize the following expression need to be found:
$\underset{w,a,b}{\mathrm{argmax}},\prod_{v\in V}p(v).$ (11)
To achieve this, it is necessary to maximize the expected value of $p(v)$:
$\underset{w,a,b}{\mathrm{argmax}},\mathbb{E}[\log p(v)].$ (12)
To do this, we take derivatives with respect to the weights and biases:
$\frac{\partial}{\partial W}\log p(v),$ (13)
$\frac{\partial}{\partial a}\log p(v),$ (14)
$\frac{\partial}{\partial b}\log p(v),$ (15)
The calculation of the derivative in Eq. 13 can be computed as:
$\begin{split}\frac{\partial\log p(v)}{\partial W}=&\left[\frac{\partial}{\partial W}(-E(v,h))|v=v_{n}\right]-\\ &\left[\frac{\partial}{\partial W}(-E(v,h))\right]\\ &=\langle v_{i}h_{j}\rangle_{data}-\langle v_{i}h_{j}\rangle_{model}.\end{split}$ (16)
In this equation, $\mathbb{E}$ represents the expectation, and $\langle\cdot\rangle_{data}$ and $\langle\cdot\rangle_{model}$ denote expectations calculated under the data and model distributions, respectively.
This results in a straightforward learning rule for performing stochastic steepest ascent in the log probability of the training data, as given by Eq. 17:
$\Delta w_{ij}=\epsilon(\langle v_{i}h_{j}\rangle_{data}-\langle v_{i}h_{j}\rangle_{model}).$ (17)
Because there are no direct connections between hidden units in an RBM, it is not necessary to know all the system configurations but rather the probabilities that connect a visible state to each of the nodes in the hidden state. Therefore, obtaining a sample of $\langle v_{i}h_{j}\rangle_{data}$ is very easy. Given a randomly selected training input, $v$, the binary state, $h_{j}$, of each hidden unit, $j$, is set to 1 with the following probability [24]:
$p(h_{j}=1|v)=\sigma(\beta(b_{j}+\sum_{i}v_{i}w_{ij})).$ (18)
where $\sigma(x)$ is the logistic function and $\beta$ is the inverse of temperature ($1/t$) and is used to control the noise in the sampling, this means that when $t\rightarrow\infty$ all the visible probabilities given $h_{j}=0$ or $1$ will be equal to one half.
Similarly, as there are no connections between the visible nodes, we can obtain a sample of $\langle v_{i}h_{j}\rangle_{data}$ given a hidden vector by setting each visible unit $v_{i}$ to 1 with the probability given by Eq. 19:
$p(v_{i}=1|h)=\sigma(\beta(a_{i}+\sum_{j}h_{j}w_{ij})).$ (19)
The problem arises when trying to obtain the expectation under the model distribution $\langle v_{i}h_{j}\rangle_{model}$ since calculating the expected value over the entire model would require knowing all possible configurations between visible and hidden states in order to obtain the probability of each one and compute the expected value [24]. This problem quickly becomes intractable [22] . For example, in an RBM with 100 nodes in the visible layer and 200 nodes in a hidden layer, there would be $2^{300}$ possible configurations, a number greater than the number of atoms in the universe ($10^{78}-10^{82}$ atoms in the universe). Storing all that information is impossible, even for a relatively simple network like the one described above.
This problem can be solved by starting at any random state of the visible units and performing alternating Gibbs sampling for a very long time. Gibbs sampling is a Markov chain Monte Carlo algorithm where each iteration consists of updating all the hidden units in parallel using Eq. 18, followed by updating all the visible units in parallel using Eq. 19.
A much faster learning procedure called Contrastive Divergence was proposed by Hinton [23]. It starts by setting the states of the visible units to a training vector. Then, the binary states of the hidden units are all computed in parallel using Eq. 18. Once the binary states have been determined for the hidden units, a “reconstruction” is produced by setting each $v_{i}$ to 1 with a probability given by Eq. 19. The change in a weight is then given by:
$\Delta w_{ij}=\epsilon(\langle v_{i}h_{j}\rangle_{data}-\langle v_{i}h_{j}\rangle_{recon}).$ (20)
In this equation, $\epsilon$ represents a learning rate, and $\langle\cdot\rangle_{recon}$ denotes expectations calculated using the reconstruction data.
RBMs typically learn better models if more steps of alternating Gibbs sampling are used before collecting the statistics for the second term in the learning rule in Eq. 20 , which will be called the negative statistics [24].
Contrastive Divergence is an algorithm to calculate maximum likelihood in order to estimate the slope of the graphical representation showing the relationship between the weights and its errors in a time complexity of each step of $\mathcal{O}(1)$.
Without Contrastive Divergence the hidden nodes in the restricted Boltzmann machine can never learn to activate properly, this method is also used in the cases where the direct probability can’t be evaluated, also, this method is one of the fastest methods to measure the log partition function.
While Contrastive Divergence facilitates efficient learning in RBMs, their true power lies in their representational capacity, enabling them to capture a wide range of distributions. Increasing the number of hidden units enhances performance by improving training log-likelihood and reducing the KL divergence between data and model distributions [25]. Research has shown that an RBM with $k+$1 hidden units can closely approximate any distribution over $\{0,1\}^{n}$, where $k$ denotes the number of input vectors with non-zero probability. Furthermore, it has been demonstrated that an RBM with $12^{n}-$ 1 hidden units can approximate any distribution [26]. Refinements in these findings suggest that RBMs with $\alpha 2^{n}$-1 hidden units (where $\alpha<1$) are sufficient for approximating any distribution [27]. This versatility makes RBMs particularly promising for addressing challenges such as Full Configuration Interaction (FCI) in quantum mechanics, where the curse of dimensionality severely limits computational tractability for complex systems. By leveraging their ability to capture complex distributions efficiently, RBMs offer a potential solution to overcoming the computational challenges posed by FCI.
Here’s how RBMs can bridge this gap:
  • 1.
    Learning the Configuration Space: An RBM can be trained on a dataset of electronic configurations relevant to the system under study. By learning the underlying probability distribution of these configurations, the RBM can efficiently sample the most important configurations for the FCI calculation.
  • 2.
    Selective Sampling vs. Random Exploration: Traditional montecarlo sampling based methods for FCI might explore the configuration space randomly, potentially wasting computational resources on irrelevant configurations. The RBM, on the other hand, prioritizes generating new configurations, giving higher likelyhoood to those that contribute more to the solution. This selective approach significantly reduces the computational cost of FCI.
The next section will explore how RBMs can be integrated with Configuration Interaction calculations to achieve faster convergence and enable the study of electronic structures in complex systems that were previously intractable.
2.4 Proposed Model
Algorithm 1 Generative Model with Determinant Pruning
  Define constants:
[…]
  $\text{Pthreshold}\in\mathbb{R}$
  $\text{Sthreshold}\in\mathbb{R}$
  $\text{epochs}\in\mathbb{N}$
  Initialize variables:
  $E\leftarrow 0$Initialize energy}
  $\text{PDetsList}\leftarrow\emptyset$initial pruned determinants list}
[…]
  for iteration $=1$ to max_iterations do
     for epoch $=1$ to epochs do
        Train RBM (dets)
     end for
[…]
     $\mathbf{D}^{\prime}\leftarrow\text{$\mathbf{D}^{\prime}$}+\text{dets}$
     $\mathbf{D}^{\prime}\leftarrow\text{Unique($\mathbf{D}^{\prime}$)}$
[…]
     $\mathbf{C}\leftarrow\text{Diagonalization($\mathbf{D}^{\prime}$)}$
[…]
     $\text{PDetsList}\leftarrow\text{PDetsList}\cup\{\mathbf{D}^{\prime}[i]\mid\mathbf{C}[i]\leq\text{Pthreshold}\}$
     $E_{\text{new}}\leftarrow\text{ComputeEnergy}(\text{dets})$
     if $|E_{\text{new}}-E|<\text{Sthreshold}$ then
        break
     end if
     $E\leftarrow E_{\text{new}}$
  end for
Our approach aims to solve the Schrödinger equation for the many-body Hamiltonian by integrating a novel data-driven method inspired by Full Configuration Quantum Monte Carlo (FCIQMC) technique but guided by a neural network RBM. Instead of stochastic exploration, our model seeks to learn the distribution of determinants that significantly contribute to the wave function. Our approach allows for a more guided exploration in the Hilbert space, reducing the number of determinants needed when diagonalizing the Hamiltonian, leading to faster and more efficient convergence to the ground state at each iteration and offering a refined approximation compared to traditional methods like Hartree-Fock or CISD.
2.4.1 Algorithm Overview
Similar to Monte Carlo methods, our model begins with an initial ansatz for the wave function, typically obtained from a CISD calculation performed using the Quantum Package software [28]. The initial set of determinants from this calculation is used to train the neural network, aiming to learn a distribution across the configuration space. New determinants are then sampled from this learned distribution using multithreading until we have a sufficient number of unique determinants that meet the criteria of single or double substitution.
These new determinants are combined with the existing training set, ensuring duplicates are removed by mapping vectors to decimals for vectorized calculation. If it is the first iteration, we create a list to store determinants that will be pruned after diagonalization. In subsequent iterations, a pruning mechanism is implemented to exclude previously pruned determinants from the generated set, preventing computational resources from being wasted on determinants which minimal contribution has been established by solving the Hamiltonian, therefore avoiding re-diagonalization.
The Hamiltonian matrix is generated with the remaining determinants and then diagonalized to obtain the coefficients. Determinants with coefficients above the pruning threshold are retained, while those below the threshold are added to the taboo list for efficient management in subsequent iterations. The new energy of the system is calculated, and if the energy difference between iterations is below the convergence threshold, the algorithm stops. Otherwise, the energy is updated for the next iteration, and the process continues with the newly generated determinants as the new ansatz for training the RBM, avoiding the need to recalculate the CISD using Quantum Package.
This iterative process is formally summarized in Algorithm 1.
2.4.2 Key Features and Innovations
Generative Model
Our approach distinguishes itself from some previously proposed Restricted Boltzmann Machine (RBM) models in several key aspects. In each iteration of our model, we utilize the determinants that emerged in the previous iteration according to their coefficients obtained through diagonalization. These same determinants are used to train the model in the subsequent iteration. This approach differs from previously proposed models, such as Carleo’s model [9], who did not use RBMs as generative models. In contrast, Basile and co-workers [10] employed a dataset of randomly sampled CISD determinants to train their model.
Handling of Pruned Determinants
A notable improvement of our model from Herzog’s approach [10] is in the handling of pruned determinants. Given that our model is generative, it may regenerate determinants in subsequent iterations that were previously diagonalized and discarded due to their negligible contribution to the wave function. To address this, we store the pruned determinants in each iteration in the beforementioned taboo list, to make sure they won’t be added again in the composition of the hamiltonian matrix of each iteration. After the generative process in each iteration, we compare the newly generated determinants with the list of pruned determinants and eliminate those that had been previously diagonalized. This step prevents re-diagonalizing determinants that were previously found unimportant, thereby saving computational time during new diagonalizations. This process has proven highly effective, significantly reducing diagonalization times.
Binary-to-Decimal Conversion
During the generative part of our process, we ensured that there were no duplicate determinants among those generated. To achieve this, we used a binary-to-decimal conversion, which allowed us to map binary vectors to decimals. This method aligned with the work of Norm M. Tubman [29] regarding the significant impact of algorithm optimization on computational time reduction, allowed us to have an algorithm a more efficient algorithm by utilizing vectorized functions that can be applied to integer arrays, facilitating faster comparison between elements. This technique helped us maintain the uniqueness of the determinants and improve the overall efficiency of the algorithm.
Physics Constraints
Our model adheres to physics constraints, ensuring the integrity of closed and symmetric systems. Specifically, we enforce conservation of the same number of alpha (spin 1/2) and beta electrons (spin -1/2). During the determinant generation process, maintaining spin symmetry ensures an equal distribution of spin-up and spin-down electrons. This adherence to physical principles is crucial as we operate within specific rules governing our models.
To achieve this, we employ a non-uniform finite distribution sampling algorithm, known as the tower sampling method [30], which has demonstrated its effectiveness in similar contexts [10]. The probability distribution utilized in the tower sampling method is derived from the distribution learned by the RBM. The implementation of this method is presented in Algorithm 2, specifically for determining the Molecular Orbitals occupied by the alpha electrons.
Algorithm 2 Tower sampling to find Occupied Molecular Orbitals for $\alpha$ electrons
  Inputs:
  $\hskip 8.5359ptN_{\alpha}=N_{e}/2$ (Numbero of $\alpha$)
  $\hskip 8.5359pt\{\pi_{1},\ldots,\pi_{N_{\alpha}}\}$ (Probabilities from RBM)
  $\Pi_{0}\leftarrow 0$
  for $l=1,\ldots,N_{\alpha}$ do
     $\Pi_{l}\leftarrow\Pi_{l-1}+\pi_{l}$ (cumsum)
  end for
  $\Upsilon\leftarrow\text{ran}(0,\Pi_{N_{\alpha}})$
  Find $k$ such that $\Pi_{k-1}<\Upsilon<\Pi_{N_{\alpha}}$
  Output: $k$ (MO occupied by a $\alpha$)
Slater’s rules
Determinants that differ by more than one or two permutations (singlets or doublets) yield elements of 0 in the Hamiltonian matrix elements, thus providing no contribution to the final wave function, in accordance with Slater’s rules.
2.4.3 Efficiency and Convergence
reach similar or even better results as current state of the art methods, with up to two orders of magnitude less determinants. This efficiency is largely attributed to storing determinants from previous iterations, avoiding constant re-diagonalization of low-contribution determinants and most importantly, increasing the likelihood of finding significant determinants by guiding the RBM. The iterative training process prioritizes determinants that contributed the most in previous iterations, enhancing the model’s ability to capture essential features of the configuration space. This strategy helps maintain the time complexity of the Davidson algorithm used to solve the Hamiltonian matrix, at $O(n^{2})$, or at worst $O(n^{3})$, preventing rapid growth in computational cost and reducing overall computation time. Our RBM-based approach can recover up to 99.99% of the correlation energy using three to four orders of magnitude fewer determinants compared to full CI calculations, as illustrated in Table 2.
2.4.4 Understanding the Learned Distribution
Part of our investigation focuses on understanding the distribution learned by the model, aiming to comprehend its effectiveness and the physics it captures. This deeper understanding helps validate the model and its adherence to physical principles.
2.4.5 Convergence Criterion
Our approach integrates a convergence criterion to ensure termination when the energy change between iterations falls below a specified threshold. This criterion is based on the observation that with each iteration, the number of determinants sampled from the configuration space increases, leading to progressively better approximations. This ensures computational resources are used effectively, avoiding unnecessary iterations once the solution has sufficiently converged.
2.4.6 Implications of Pruning and Determinant Management
Implications of Pruning and Determinant Management: Several assumptions and approaches in our algorithm influence the final energy approximation. For instance, using a threshold to prune determinants with coefficients below it may lead to an inaccurate final energy estimation. However, the threshold is deliberately set well below the desired energy order, ensuring that even with significant determinant elimination, the energy approximation remains within an acceptable range.
Moreover, the selective elimination of determinants based on their coefficients can bias the energy approximation. Eliminating only determinants with positive coefficients might continuously decrease the energy, potentially yielding a more negative energy than the actual ground state. Conversely, eliminating only negative coefficients could inflate the energy, deviating from the true ground state. To mitigate these biases, we use the squared absolute value of the coefficient ($|c|^{2}$) as the pruning criterion. This approach ensures a balanced removal of determinants based on their contribution magnitude rather than sign, thereby reducing previous errors.
Furthermore, our practice of storing and excluding previously pruned determinants during subsequent iterations does not guarantee that a previously pruned determinant won’t be relevant in future iterations with new determinants. While theoretically possible, it’s improbable due to the low threshold used and the tendency of subsequent determinants to represent higher-order excitations. Higher-order excitations typically exhibit lower probabilities or coefficients, making it unlikely for a previously pruned determinant to become significant in future iterations.
Considering these implications, our model’s approximation tends to improve with increased exploration of the Hilbert space, reflecting a closer approximation to the desired energy state as iterations progress.
3 Results and Discussion
In this section, we present the results obtained using the proposed model for the molecules $H_{2}O$, $N_{2}$, and $C_{2}$ in the 6-31G and cc-pVDZ basis sets. The energy convergence across iterations for these molecules is depicted in Figure 2. It is evident how our model progressively converges towards FCI energy levels, demonstrating the effectiveness of our approach.
Figure 2: Energy convergence across iterations for molecules of $H_{2}O$, $N_{2}$, and $C_{2}$ in cc-pVDZ and 631G basis.
To provide a comprehensive comparison, Table 2 shows the energy approximations obtained using CISD, FCI, and our proposed model across different molecules. The results indicate that our model’s energy values approach to FCI levels after a certain number of iterations. Focusing on the $N_{2}$ molecule, our model achieves correlation energies of 99.99% with 378,000 determinants, which is a significant improvement compared to the work of [10]. In their study, they obtained 99.979% of the correlation energy for $N_{2}$ in the cc-pVDZ basis using 19 million determinants, and only 98.120% with 156,000 determinants. Moreover, when comparing the number of determinants used in our solution with the total number listed in Table 1 for FCI, it is evident that our approach achieves similar levels of accuracy with 3 to 4 orders of magnitude fewer determinants
Molecule Basis CISD energy FCI energy Our energy % of Correlational Energy our Num Dets Iterations
$H_{2}O$ cc-pVDZ -76.2320 -76.24195 -76.24219 99.999 182655 6
6-31G -76.1139 -76.12237 -76.12360 99.998 27636 6
$C_{2}$ cc-pVDZ -75.6669 -75.72984 -75.73522 99.992 231026 5
6-31G -75.5846 -75.64418 -75.65191 99.989 82413 4
$N_{2}$ cc-pVDZ -109.2438 -109.27834 -109.28322 99.995 378477 8
6-31G -109.0759 -109.10842 -109.10886 99.999 200983 6
Table 2: Comparison of Energy Approximations and Computational Effort: A Comparative Analysis between CISD, FCI, and Our Proposed Model, Including the Number of Determinants and Iterations Required for Energy Convergence.
Our method also exhibits remarkable computational efficiency, even when operating with a large number of determinants. Figure 3 illustrates that the computational time increases quadratically with the number of determinants for molecules $H_{2}O$, $N_{2}$, and $C_{2}$ using the cc-pVDZ basis set. This quadratic behavior is expected due to the Davidson algorithm’s time complexity, which typically operates at $O(n^{2})$. However, it is important to note that the computational time remains manageable even as the number of determinants increases substantially. This highlights our model’s capability to handle large-scale simulations without exploiting computational complexity.
Figure 3: Computational time required in function of the number of determinants in each iteration.
The efficient use of computational resources can be attributed to several factors inherent in our approach. Firstly, by storing determinants from previous iterations and avoiding constant re-diagonalization of those with low contributions to the wave function, we focus computational efforts on the most impactful determinants. Moreover, our iterative training process prioritizes determinants that contributed the most in previous iterations.
Through a grid search that varied the number of epochs, temperature ($1/\beta$ in 18 and Eq. 19), batch size, and number of Gibbs sampling steps, it was found that the models consistently performed best when a temperature of $11/\beta=$ 1 is used, signifying a focused effort to learn the distribution entirely. Conversely, models with higher temperatures ($51/\beta=$ 5 and $301/\beta=$ 30), indicative of more stochastic behavior, exhibited inferior performance across nearly all cases, with temperature $301/\beta=$ 30 (representing random sampling) showing the least proximity to the FCI energy, as observed in Figure 4 for the $H_{2}O$ molecule in the cc-pVDZ basis. Additionally, it was observed that in all temperature settings, fewer epochs, small batch sizes, large numbers of Gibbs samplings, and low learning rates were associated with better performance compared to using more epochs. This suggests that the model seeks to approximate the general distribution at each iteration but avoids fully internalizing the current training distribution, enabling it to sample determinants out of the training distribution.
Figure 4: Hyperparameter search results (only 4 iterations) showing energy (x-axis) versus epochs (y-axis) for $H_{2}O$ molecule in the cc-pVDZ basis for different temperatures $51/\beta=1,$ 5 and $30$. Models with lower temperatures and fewer epochs performed best, approaching FCI energy more closely.
Analyzing the occurrence occupation in each Molecular Orbital (MO) for the water molecule in the 6-31G basis, Figure 5 shows a transition from a Fermi-Dirac-like distribution to one still predominantly concentrated on the lowest energy levels (first MOs for alpha and beta electrons), although with a distribution that seems oscillatory. This graph demonstrates how often each of the 26 MOs (13 alpha and 13 beta) was indicated as occupied during the generation of determinants in each iteration, providing insight into the frequency with which each orbital was utilized as occupied for water in the 6-31G basis.
These results may explain why the model with a high temperature performs poorly in Figure 4. This can be attributed to the fact that RBMs with completely random training moves have difficulty sampling multimodal distributions ergodically [31]. At the same time, this explain why the models with a temperature of $11/\beta=$ 1 works better with fewer epochs. Multimodal distributions are notoriously hard to sample because simulations often get trapped in one or more metaestable energy minima for extended periods, failing to explore the entire configuration space. For this reason, fewer epochs are needed to avoid overfitting and reduce the risk of getting stuck in local minima.
Figure 5: Comparison between Frequency occupation in each Molecular Orbital(blue bars) and the frequency weighted by the coefficients obtained in diagonalization (green bars) for $H_{2}O$ in 6-31G basis for each iteration.
A similar behavior is observed in the frequency of each MO weighted by the coefficients obtained from each diagonalization ($c_{i}$ in Eq.5), as shown in Figure 5 with green bars for $H_{2}O$ in the 6-31G basis. These coefficients represent the contribution of each determinant or its probability ($|c_{i}|^{2}$), allowing us to infer the occupation based on the same probability. Interestingly, the model appears to learn a distribution that might be associated with a relationship between the frequencies with which an orbital is occupied and the underlying distribution of the determinants. Despite starting from random data, the model consistently learns the same distribution when repeated multiple times. This pattern shows certain differences between different molecules, but notable similarities between the same molecule in different bases. This is evident in the patterns persisting across similar molecules in different bases, as shown in Figure 6, while varying for molecules of different compounds.
Figure 6: Final frequency occupation for each MO for for molecules (columns ) of $H_{2}O$, $N_{2}$, and $C_{2}$ in cc-pVDZ and 631G basis (rows).
We ensured that the RBM learned a probability distribution function by comparing two experiments for the same molecule. In one experiment, we randomly reinitialized the weights in each iteration, while in the other experiment, we did not reinitialize them. Remarkably, in both cases, the distributions obtained in each iteration exhibited the same behavior (Figure 7). This indicates that the RBM, at each iteration, is attempting to discern the distribution of Molecular Orbitals (MOs).
Figure 7: Comparison between frequency occupation in each Molecular Orbital reinitialized and not reinitialized in each iteration for $H_{2}O$ in 631G basis for each iteration.
In Figure 8, we observe a very similar energy convergence towards the FCI Energy for both cases. This similarity provides further evidence that the oscillatory pattern is not merely a feedback artifact generated by the RBM, but rather a learned feature of the system itself.
The information of Figures 6 and 7 indicates that the model tends to fit the same distribution regardless of the molecular basis or weight initialization. This suggests ergodic sampling, as the distribution consistently adjusts with each iteration. To verify that our sampling is ergodic and does not lead to non-convergence Markov chains, we can use the fact that the RBM is not trained using gradient descent, as in other neural networks, but through Gibbs sampling, which is a specific instance of a Markov Chain Monte Carlo (MCMC) method [32]. By treating the RBM in this way, we can apply the Gelman-Rubin statistic to assess convergence. Consider $J$ Markov chains, each initialized with different random values. For each chain $i$ of $L$ steps, the mean value of the chain is given by:
$\bar{x}_{j}=\frac{1}{L}\sum_{i=1}^{L}x_{i}^{(j)}$ (21)
The mean of the means across all chains is:
$\bar{\bar{x}}_{j}=\frac{1}{J}\sum_{j=1}^{J}\bar{x}_{j}$ (22)
The variance of the means of the chains, also known as the between-chain variance $B$, is defined as:
$B=\frac{L}{J-1}\sum_{j=1}^{J}\left(\bar{x}_{j}-\bar{\bar{x}}_{j}\right)^{2}$ (23)
The average variance of the individual chains across all chains, known as the within-chain variance $W$, is:
$W=\frac{1}{J}\sum_{j=1}^{J}\left(\frac{1}{L-1}\sum_{i=1}^{L}(x_{i}^{(j)}-\bar{x}_{j})^{2}\right)$ (24)
The Gelman-Rubin statistic $R$ is then defined as [33]:
$R=\sqrt{\frac{\frac{L-1}{L}W+\frac{B}{L}}{W}}$ (25)
As $L$ tends to infinity and $B$ tends to zero, $R$ tends to 1, indicating convergence.
In our case, since the different runs of the algorithm converge to similar distributions after a certain number of iterations, and the energy values are consistent across these runs, it is as if we were running $J$ simulations where the variance of the means of the chains is very low. For instance, the variance of the normalized distributions between the reinitialized and non-reinitialized weights is 3 ×$\times 10^{-4}$, which corresponds to the value of $B$ in the Gelman-Rubin equation. This low variance implies that $B$ is close to zero, leading to an estimate of the Gelman-Rubin statistic $R$ that is very close to 1. This result suggests that our sampling process is ergodic and not subject to non-convergence issues.
Figure 8: Energy convergence for the experiments with reinitialized and not reinitialized weight on each iteration.
Another analysis confirming the RBM’s accurate learning of a distribution can be seen in Figure 9. Here, we adapt Gaussian distributions to the weighted occupancy frequency graphs in Figure 6, aiming to extract valuable insights. The Gaussian distributions exhibit similar behaviors, with their means (centering of the Gaussian) positioned at the same MOs and displaying comparable behaviors across different molecular bases. However, for the cc-pVDZ basis configuration, the model can sample MOs with higher energies than the 6-31G basis.
Figure 9: Gaussian distributions fitted to the weighted occupancy frequency graphs for different molecules and bases, illustrating similar behaviors. The Gaussian centers are conserved between bases for the same molecule, although the cc-pVDZ basis covers a larger MO space. Where they overlap, the patterns are comparable
Analyzing the behaviors of the envelopes function of Figure 9, a similarity to the radial probability depicted in Figure 10 emerges. This illustration shows a surface probability $\psi^{2}r^{2}$ against $r$, demonstrating that electrons occupying higher energy orbitals are more likely to be farther from the nucleus. These unique yet consistent oscillations across different bases may signify a connection to the molecule’s electronic configuration.
Figure 10: Radial Distribution Function (RDF) for a hydrogen atom, showing radial probability $\psi^{2}r^{2}$ against distance $r$ from the nucleus in units of $a_{0}$. The graph displays the 1s, 2s, 2p, and 3s orbitals.
We also observe a similar behavior with the obtained Radial Distribution Function (RDF) in experiments, rigid models, and Monte Carlo simulations for $H_{2}O$ [34] [35], $C_{2}$ [36] [37] and $N_{2}$ [38] For instance, in our results for a molecule of $H_{2}O$, the lowest molecular orbitals exhibit two prominent peaks (see Figure 9), similar to patterns observed in the Radial Distribution Function (RDF) experiments at low radial distances. These peaks, as indicated by previous studies, are associated with the O-H bond structure.[35], [34].
While a complete correlation between our results and experimental RDFs cannot be established due to the inherent differences between atomic and molecular electron distributions, the observed similarities are notable. In atoms, electrons are distributed around a single nucleus in a predictable manner, whereas in molecules, neighboring atoms influence this distribution. These influences are three-dimensional and depend on the specific atomic positions. Although rules exist for combining atomic orbitals leading to constructive or destructive interference, these interactions manifest in complex three-dimensional spaces. Therefore, the applicability of a radial density function (RDF) may be limited, especially when compared across different molecules.
Finally, an important analysis is observed in Figure 11, where we compare the distribution learned by the RBM with the normalized orbital energies for each molecule in the cc-pVDZ basis. The orbital energies were computed using the software Q-chem [39]. Remarkably, the phenomenon of having two or four prominent energy peaks in the lowest states is observed in the distribution learned by the models (Fig 9). This suggests that the model might be learning relevant information about orbital energy, particularly for those with the lowest energy levels.
Figure 11: Comparison of the frequency distribution learned by the RBM with the normalized orbital energies for each molecule in the cc-pVDZ basis
4 Conclusion
Our work presents a novel approach to enhance Full Configuration Interaction Quantum Monte Carlo (FCIQMC) by integrating a Restricted Boltzmann Machine (RBM) with an novel iterative training strategy, based on the prioritizing the determinants that most contribute to the wave function, and the use of a taboo list to learn the distribution of key determinants, thus avoiding sampling the entire configuration space.
This strategy offers several advantages. Firstly, by focusing on determinants with a higher contribution, our method requires sampling a smaller number of such determinants, compared to FCI, leading to faster and more efficient convergence towards the ground state energy, significantly reducing computational cost.
Secondly, unlike traditional Monte Carlo approaches that randomly sample the entire configuration space, our method utilizes the RBM to guide the exploration towards more relevant regions. This guided exploration leads to improved convergence.
Thirdly, the model adheres to physical constraints such as a constant number of electrons and spin symmetry, ensuring physically meaningful results.
The effectiveness of this approach is demonstrated through various observations. The model exhibits consistent energy convergence towards FCI energy levels for various molecules and basis sets, showcasing its efficiency (Figure 2). Analysis of hyperparameter optimization suggests the model prioritizes learning the general distribution across iterations, enabling exploration beyond the current state and potentially contributing to faster convergence.
Furthermore, the inherent simplicity of the RBM network allows us to provide explainability by the analysis of frequency occupation in each Molecular Orbital (MO) reveals a transition towards a distribution focused on the lowest energy levels (Figures 5 and 6). This behavior resembles the contribution probabilities of determinants to the wave function and holds true across different molecules. We confirmed the RBM learns a probability distribution function by demonstrating consistent distribution behavior regardless of weight reinitialization during training (Figure 7).
In addition, the model-generated envelope function exhibits similarities with established radial probability densities (Figure 10), potentially reflecting a connection to the molecule’s electronic configuration. Additionally, the RBM distribution captures relevant information about orbital energies, particularly for the lowest energy levels (Figure 11).
These findings collectively suggest that our RBM-based approach holds significant promise for enhancing FCIQMC. It offers a more efficient and guided exploration of configuration space while capturing relevant physical properties. Further investigations are needed to explore the applicability of this method to larger and more complex molecular systems, potentially leading to breakthroughs in the study of electronic structures.
CRediT authorship contribution statement
Jorge I. Hernandez-Martinez: Conceptualization, Methodology, Software, Formal analysis, Investigation, Writing - Original Draft.
Gerardo Rodriguez-Hernandez: Resources, Supervision, Writing - Review and Editing, Guidance in the physics domain.
Andres Mendez-Vazquez: Resources, Supervision, Writing - Review and Editing, Guidance in neural networks.
Acknowledgements
The authors acknowledge the technical support of SL Juárez Osorio. This work was partially supported by CONAHCYT, Mexico
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Supplementary Information
This research stands out due to its practical implementation in Python, in contrast to previous approaches that often relied on programming languages such as Fortran. This choice not only facilitates the accessibility and interoperability of the proposed method but also allows for greater flexibility and modularity in its application and future development.
The code are available at: